OpenStudy (love_to_love_you):

Please help me, will give medal and fan!! :) Solve the equation. Check for extraneous solutions. Show all work. 9|9-8x| = 2x+3

OpenStudy (love_to_love_you):

Can y'all help?

OpenStudy (love_to_love_you):

@k_lynn

OpenStudy (anonymous):

9|9-8x|= 2x+3 81-72x=2x+3 -3+72x+72x-3 _____________ 78=78x x=1

OpenStudy (love_to_love_you):

Thank you!!